If A,B,C are the angles of triangle ABC and tanA,tanb,tanc are the roots of the equation x^​4-​3​x^​3+​3x^​2+​2​x+​5 then find the 4th root of the equation.

Grade:12

2 Answers

Lab Bhattacharjee
8 years ago
\text{ We can prove for } A+B+C=\pi, \tan A+\tan B+\tan C=\tan A\tan B\tan C\\ \text{Let this } =t \text{ and the fourth root be } y\\ \text{So, we have } t+y=\dfrac31\text{ and } ty=\dfrac51
Karthik Pasupulatei
8 years ago
use the formula alpha +beta+gama+delta =-(-3) and again uing these actors find the value